OpenStudy (dmezzullo):

I have a question. I was just looking at a problem and was gonna help someone but the other person put the work and answer first And they did it differently then me but both are still right.

OpenStudy (amistre64):

thats a statement, not a question ....

OpenStudy (dmezzullo):

Lemme post the problem

OpenStudy (amistre64):

if they "put the work and answer" then they are not conforming to the rules for this site ...

OpenStudy (dmezzullo):

Oh

OpenStudy (dmezzullo):

The problem was 300% of 18 so i did 300 x .18 and got 54

OpenStudy (dmezzullo):

Nvm im confusing myself

OpenStudy (amistre64):

1% is simply one out of a hundred if we divide 18 into 100 pieces we get 1% = 18/100 300 of those pieces amounts to 300 x 18/100 or\[\frac{300(18)}{100}=300(.18)=3(18)\]

ganeshie8 (ganeshie8):

yes all we have to do is to replace % wid 1/100

OpenStudy (dmezzullo):

Ok but i left out the 100 part, and just went straight too the answer of 54

OpenStudy (amistre64):

you just used the middle form

OpenStudy (amistre64):

they are all equal formats

OpenStudy (dmezzullo):

Is there anything wrong with doing it like that?

OpenStudy (amistre64):

tell me: is there anything worng with using say; 5+3 instead of 3+5?

OpenStudy (dmezzullo):

No sir

OpenStudy (amistre64):

is there anything wrong with using say: 6(2) instead of 2(6) ?

OpenStudy (dmezzullo):

No sir

OpenStudy (amistre64):

then there is nothing wrong with using 3(.18) or 3(18) since both are equal to 300 pieces of 18/100

OpenStudy (amistre64):

lol ... 300(.18) and 3(18) that is :)

OpenStudy (dmezzullo):

Alrighty thanks :)

OpenStudy (amistre64):

as one last consideration: 300(.18) = 3(100)(.18) = 3(18) right?